# Enhancing Speaker Verification with w2v-BERT 2.0 and Knowledge Distillation guided Structured Pruning

### v2

@@ 1 Introduction
Speaker verification (SV) aims to authenticate the identity of a speaker by analyzing the voice signal. In recent years, significant advancements in deep learning, coupled with the availability of large-scale labeled datasets [1, 2, 3, 4, 5], have led to substantial improvements in the performance of deep neural network-based SV systems [6, 7, 8].
@@
Previous studies focused on Transformer-based self-supervised PTMs for the SV task [13, 14, 15, 16]. In contrast, w2v-BERT 2.0 [12] is a self-supervised PTM built on a Conformer-based architecture, which has been demonstrated by MFA-Conformer [19] to be effective for SV and superior to the Transformer-based architecture. Furthermore, w2v-BERT 2.0 adopts a training strategy that optimizes both a contrastive loss and a masked prediction loss simultaneously, and it is trained on 4.5 million hours of unlabeled audio covering 143 languages, leading to strong performance on audio classification tasks.
@@ 2.6 Structured Pruning with Knowledge Distillation
where $u_{j}$ is drawn from a uniform distribution, $\beta$ is a temperature parameter controlling the smoothness of $s_{j}$, and $\zeta$ and $\gamma$ control the upper and lower bounds of $s_{j}$. We set $\beta=2/3$, $\zeta=-0.1$, and $\gamma=1.1$. Finally, the expected value of the L0 norm is given by:
@@ 4 Results
Table 3: Results of knowledge distillation guided structured pruning on a w2v-BERT 2.0 based SV model trained on VoxCeleb2 and VoxBlink2.
Model Sparsity Params LMFT Vox1-O EER
LoRA_Adapter_MFA 0% 580+6.2M $\times$ 0.23%

### v3

@@ 1 Introduction
Speaker verification (SV) aims to authenticate the identity of a speaker by analyzing the voice signal. In recent years, significant advancements in deep learning, coupled with large-scale labeled datasets [1, 2, 3, 4, 5], have led to substantial improvements in the performance of deep neural network-based SV systems [6, 7, 8].
@@
Previous studies focused on Transformer-based self-supervised PTMs for the SV task [13, 14, 15, 16]. In contrast, w2v-BERT 2.0 [12] is a self-supervised PTM built on a Conformer-based architecture, which has been demonstrated by MFA-Conformer [19] to be effective for SV and superior to the Transformer-based architecture. Furthermore, w2v-BERT 2.0 adopts a training strategy that optimizes both a contrastive loss and a masked prediction loss simultaneously, and trained on 4.5 million hours of unlabeled audio covering 143 languages, leading to strong performance on audio classification tasks.
@@ 2.6 Structured Pruning with Knowledge Distillation
where $u_{j}$ is drawn from a uniform distribution, $\beta$ is a temperature parameter controlling the smoothness of $s_{j}$, and $\zeta$ and $\gamma$ control the upper and lower bounds of $s_{j}$. We set $\beta=2/3$, $\gamma=-0.1$, and $\zeta=1.1$. Finally, the expected value of the L0 norm is given by:
@@ 4 Results
Table 3: Results of knowledge distillation guided structured pruning on a w2v-BERT 2.0 based SV model trained on VoxCeleb2 and VoxBlink2. MACs and FLOPs were measured on 1-s long segments.
Model Sparsity Params MACs FLOPs LMFT Vox1-O EER
LoRA_Adapter _MFA 0% 580+6.2M 28.75G 57.72G $\times$ 0.23%